User:Steven J. Koch/Notebook/Kochlab/2009/07/08/Feedback 96, QPD

From OpenWetWare

Jump to: navigation, search

Steve Koch 01:11, 9 July 2009 (EDT): Been messing around for a while now, and getting a bit tired. Clear that I still have some programming to do, especially in regards to using the piezo conveniently (e.g., how to have it automatically at the starting point, to avoid snap at the beggining of experiments, etc.).

  • The first time I looked at the trap tonight, it did not look as nice as earlier. I did not notice knocking anything: the only thing really possible would be from moving the beam stop, but I was pretty careful. Hopefully this does not indicate drift of some element. In any case, it still tweezes.
    • Andy Maloney 01:44, 9 July 2009 (EDT): I think I hit something. Grrr. Argh! Damn it all to hell....I'll fix it tomorrow. I think I hit one of those mirror mounts from Newport. I did also notice that when you are "fishing" the beam will change a bit depending on where you are in the sample.
  • Andy got everything set up really nicely. Some things that are missing:
  • More BNC cables. T-connectors
    • Andy Maloney 01:44, 9 July 2009 (EDT): If I get bored in the morning then I will make some BNC cables. The T-connectors are in the cabinet with the optics in a drawer that says "BNC Connectors".
    • Steve Koch 01:54, 9 July 2009 (EDT): It'd probably be a better use of your time to just buy some or borrow from someone in the building?
  • Need other computer for oscilloscope
    • Andy Maloney 01:44, 9 July 2009 (EDT): Doesn't that computer have 2 processors? If it does, can't you use one for the feedback and the other for the oscilloscope?
      • Possibly...but w/ two DAQ cards and two monitors anyway, it may be just as easy to use two computers

Feedback necessities

  • Modify program (possibly) to do the math for the QPD: divide X and Y by the sum (this is how analysis software is written)
  • Figure out why Power Spectrum is not working

Viscous drag test

Pin 2 on the left, pin 6?, pin 7 on the right. Laser 10.15 amps

I took a video of the bead being displace in the trap. I also took a picture of the oscilloscope screen, showing that it is easy to detect. Conditions:

  • Laser 10.15 amps
  • Using "Test Panels" for sine wave; 726 mV amplitude, 2000 update rate (2 Hz overall)

power spectrum w/ scope

Looking at FFT on scope

  • 11.39 A (0.98 Watt): rolls off like 3 boxes
  • cycle RMS 80 - 180 mV
  • 10.15 A (0.47 Watt): rolls off in like 2 boxes
  • cycle RMS 100 - 140 mV
  • 9.17 A (0.15 Watt): rolls off in 1 or less boxes
    • cycle RMS 30-80 mV
Personal tools